OpenStudy (anonymous):

Amir lent Rs. 600 to Hameed for 2 years and Rs. 150 to Aman for 4 years and received altogether from both Rs. 90 as simple interest. The rate of interest is a 4% b 5% c 10% d 12%

OpenStudy (anonymous):

Simple interest formula: I = Prt (interest = amount lent * interest rate * time) Interest from $600 for 2 years: I = 600*r*2 = 1200r Interest from $150 for 4 years: I = 150*r*4 = 600r Total interest 1200r + 600r = 90 solve for r

OpenStudy (anonymous):

thanks very much,it is 5%,thanks for helping